OpenStudy (anonymous):

Help please, I'm clueless... (2(x^2 + x + 1)^1/4 = (16x^2 + 10x - 14)^1/4

OpenStudy (anonymous):

On the left hand side you opened 2 parenthesis but only closed one. I'm assuming this is what you meant. And you are solving for x? \[2(x^2+x+1)^\frac{ 1 }{ 4 } = (16x^2+10x-14)^\frac{ 1 }{ 4 }\] If you raise both sides to the 4th power you get: \[16(x^2+x+1)=16x^2+10x-14\] Distribute the 16 on the left side. \[16x^2+16x+1 = 16x^2+10x-14\] If you subtract everything from the right hand side you get. \[6x+30=0\] Factor \[6(x+5)=0\] Divide by 6 \[x=5\]

OpenStudy (anonymous):

Thank you!